OpenStudy (anonymous):

Solve for x. x - 1/12 = 7/2 The answer is x = 43.

OpenStudy (thomas5267):

There are a few ways of doing it. I would first multiply the whole equation by 12 because I don't like fractions.

OpenStudy (anonymous):

Oh, I know how to do it, I just need my answer to be checked. My answer is 43 = x. That is correct, right?

OpenStudy (anonymous):

@thomas5267 ^

OpenStudy (thomas5267):

I am afraid that the answer is wrong. 43-1/12=515/12

OpenStudy (anonymous):

No, I am positive that it is correct.

OpenStudy (anonymous):

43 - 1 = 42 42/12 = 7/2

OpenStudy (anonymous):

The two answers are equal.

OpenStudy (thomas5267):

Oh! So it is \(\frac{x-1}{12}=\frac{7}{2}\) not \(x-\frac{1}{12}=\frac{7}{2}\)?

OpenStudy (ccswims):

in order to solve for 'x' you need to do the inverse operations...

OpenStudy (ccswims):

so instead of - 1/12 you add 1/12 to both sides

OpenStudy (anonymous):

Yeah! Sorry if it was confusing! @thomas5267

OpenStudy (thomas5267):

Then yes your answer looks right!

OpenStudy (ccswims):

OH, ok I thought that looked too easy, I get it :P

OpenStudy (anonymous):

@ccswims I actaully just multiplied 12 by 7, then I divided 84 by 2, to get 42, which is 1 less than 43. :)